# S-DMs: Skip-Step Diffusion Models

### 03 Jan 2024 v1

@@ Abstract
Diffusion models have emerged as powerful generative tools, rivaling GANs in sample quality and mirroring the likelihood scores of autoregressive models. A subset of these models, exemplified by DDIMs, exhibit an inherent asymmetry: they are trained over steps but only sample from a subset of during generation. This selective sampling approach, though optimized for speed, inadvertently misses out on vital information from the unsampled steps, leading to potential compromises in sample quality. To address this issue, we present the S-DMs, which is a new training method by using an innovative , meticulously designed to reintegrate the information omitted during the selective sampling phase. The benefits of this approach are manifold: it notably enhances sample quality, is exceptionally simple to implement, requires minimal code modifications, and is flexible enough to be compatible with various sampling algorithms. On the CIFAR10 dataset, models trained using our algorithm showed an improvement of 3.27% to 14.06% over models trained with traditional methods across various sampling algorithms (DDIMs, PNDMs, DEIS) and different numbers of sampling steps (10, 20, …, 1000). On the CELEBA dataset, the improvement ranged from 8.97% to 27.08%. Access to the code and additional resources is provided in the github.
@@ 1 Introduction
Generative models, especially deep generative models, play a foundational role in the machine learning domain((Karras et al., 2020); (Oord et al., 2016)). Architectures like Variational Autoencoders (VAEs; (Kingma & Welling, 2013)) and Autoregressive models(Van den Oord et al., 2016);(Brown et al., 2020); (Salimans et al., 2017), Generative Adversarial Networks (GANs; (Goodfellow et al., 2014); (Yu et al., 2017); (Hjelm et al., 2017);(Fedus et al., 2018)), and Restricted Boltzmann Machines (RBMs; (Hinton, 2012)) have been at the forefront. VAEs, while providing a structured probabilistic framework, occasionally yield blurry samples. GANs, acclaimed for their prowess in generating high-resolution images, can face training instabilities (Adler & Lunz, 2018); (Gulrajani et al., 2017); (Karras et al., 2019). RBMs, though seminal, find themselves overshadowed by more recent architectures in scalability and performance. Against this backdrop, diffusion models, Denoising diffusion probabilistic models(DDPMs; (Ho et al., 2020); (Sohl-Dickstein et al., 2015);(Song et al., 2020b)), have emerged as a compelling alternative, exhibiting unmatched capabilities in generating superior samples in diverse domains, from image synthesis to molecule design (Bengio et al., 2014).
@@
Driven by these observations, our research proposes a method of integrating skip-step sampling as an optimization target into the training process. The model trained in this way can consider and adapt to the information lost during the sampling period due to skip-step sampling. Our model not only maintains efficient generation when using other accelerated sampling algorithms but also improves the quality of the generated samples, thereby achieving a balance of efficiency and performance. Specifically, during the training process, the original loss function is retained, and at the same time, a skip-step loss is introduced, measuring the difference between the model’s current step prediction and the skip-step result. This skip-step loss is combined with the original loss using a weighted mechanism. Figure 1 shows an overview of the diffusion models reverse process and details the focus of our model.
@@ 3.2 Training with Skip-Step Loss
We minimize the squared difference between our desired optimization target and the original target to minimize their discrepancy, thereby achieving the desired outcome,
After performing calculations using the formula, we can obtain our final loss function ,
Due to the modification of some weight values, it is essential to introduce an appropriate weight in the subsequent sections to ensure the system functions properly. Currently, the value of this loss function is stable and does not lead to collapse (see the experimental in section 4).
@@ 4 Experiments
In consideration of computational resources, our experiments utilized the CIFAR10 dataset with a resolution of 32×32 and the CelebA dataset with a resolution of 64×64. For the training setup, we adopted the same architecture (He et al., 2016); (Ronneberger et al., 2015); (Kingma & Ba, 2014) as provided in the official DDIMs repository. We also ensured that all parameters were kept consistent, guaranteeing the reproducibility of our experiments. For hardware, we trained both datasets on two NVIDIA A100 GPUs, with 600K steps for CIFAR10 and 400K steps for CelebA. Model performance was evaluated based on the FID (Heusel et al., 2017); (Jolicoeur-Martineau et al., 2020). Specifically, Our evaluation method followed the DDIMs repository, wherein we sampled 50,000 images and computed the FID against real images. To further guarantee reproducibility, we fixed the random seed in our experiments, ensuring that all results are replicable.
@@ 4.1 Sample Quality and Ablation Experiment
In Figures 6 and 6, we display samples from the CIFAR10 and CelebA datasets generated by models with identical numbers of sampling steps and the same sampling algorithm. These figures reveal that, under the same sampling conditions, the samples generated by S-DMs are of higher quality. Additionally, as the number of sampling steps increases, so does the quality of the samples. For instance, in CIFAR10, the images of boats and cars in the second and fourth columns demonstrate that S-DMs can produce highly accurate images in just 10 steps, whereas DDIMs yield blurrier images and require more steps for high-quality outcomes. Similarly, in CelebA, the third column presents images of a male subject, where S-DMs generate a clear hat, in contrast to the blurrier hat produced by DDIMs. Moreover, the detail in images generated by PNDMs is less pronounced compared to those by S-DMs. This difference is significant, highlighting the substantial improvement in sample quality achieved by S-DMs with the same number of sampling steps.
@@ 4.2 Interpolation and Generation Consistency
Given that the S-DMs is based on the deterministic generation process of DDIMs(Song et al., 2020a), PNDMs(Liu et al., 2022a) and DEIS(Zhang & Chen, 2022), it also exhibits the semantic interpolation effects observed in implicit models(Mohamed & Lakshminarayanan, 2016), such as GANs. In Figure 7, we display the interpolation results of the S-DMs under different skip-step settings. From the figure, it can be discerned that simple interpolation in can lead to semantically meaningful interpolations between two samples. Moreover, the generated samples at demonstrate superior quality and finer details, as exemplified by the 5th to 7th images, where the light and shadow effects on the faces are also effectively reproduced. In contrast, at , the model closely resembles the original DDIMs, resulting in a loss of sample detail. Additionally, the figure illustrates that even models trained with different skip-step settings, when conditioned on the same encoding, still produce fairly consistent samples.
@@ 5 Related Work
Denoising Diffusion Probabilistic Models (DDPMs; (Ho et al., 2020)) and Noise Conditional Score Networks (NCSNs; (Song & Ermon, 2019)) are notable for their sample quality, comparable to GANs. While DDPMs optimize a variational lower bound, NCSNs target score matching over a Parzen density estimator. Both models employ a denoising autoencoder across noise levels and use Langevin dynamics for sampling. The shared approach requires multiple iterations for optimal sample quality. Recent advancements aim to decrease DDPMs’ inference steps through dynamic SDE solvers and programming algorithms. However, challenges like the disparity between log-likelihood reduction and FID (Heusel et al., 2017), (Szegedy et al., 2016) remain in some models.
@@ 6 Conclusion and Future Work
Our findings pave a new direction for future diffusion model research. By investigating the asymmetry between the training of diffusion models and accelerated sampling algorithms, we can ensure that the trained models are better suited to accelerated sampling algorithms. This ensures that even with algorithms using fewer sampling steps, the model can still generate high-quality outputs. Perhaps this offers an effective solution to the challenges diffusion models face in balancing sampling speed and sample quality. In the future, we will continue to explore how to better integrate leap step information to ensure that trained models are more consistent with accelerated sampling algorithms, thus improving sample quality.

### 23 Jan 2024 v2

@@ Abstract
Diffusion models have emerged as powerful generative tools, rivaling GANs in sample quality and mirroring the likelihood scores of autoregressive models. A subset of these models, exemplified by DDIMs, exhibit an inherent asymmetry: they are trained over steps but only sample from a subset of during generation. This selective sampling approach, though optimized for speed, inadvertently misses out on vital information from the unsampled steps, leading to potential compromises in sample quality. To address this issue, we present the S-DMs, which is a new training method by using an innovative , meticulously designed to reintegrate the information omitted during the selective sampling phase. The benefits of this approach are manifold: it notably enhances sample quality, is exceptionally simple to implement, requires minimal code modifications, and is flexible enough to be compatible with various sampling algorithms. On the CIFAR10 dataset, models trained using our algorithm showed an improvement of 3.27% to 14.06% over models trained with traditional methods across various sampling algorithms (DDIMs, PNDMs, DEIS) and different numbers of sampling steps (10, 20, …, 1000). On the CELEBA dataset, the improvement ranged from 8.97% to 27.08%.
@@ 1 Introduction
Generative models, especially deep generative models, play a foundational role in the machine learning domain (Karras et al., 2020); (Oord et al., 2016). Architectures like Variational Autoencoders (VAEs; (Kingma & Welling, 2013)) and Autoregressive models(Van den Oord et al., 2016);(Brown et al., 2020); (Salimans et al., 2017), Generative Adversarial Networks (GANs; (Goodfellow et al., 2014); (Yu et al., 2017); (Hjelm et al., 2017);(Fedus et al., 2018)), and Restricted Boltzmann Machines (RBMs; (Hinton, 2012)) have been at the forefront. VAEs, while providing a structured probabilistic framework, occasionally yield blurry samples. GANs, acclaimed for their prowess in generating high-resolution images, can face training instabilities (Adler & Lunz, 2018); (Gulrajani et al., 2017); (Karras et al., 2019). RBMs, though seminal, find themselves overshadowed by more recent architectures in scalability and performance. Against this backdrop, diffusion models, Denoising diffusion probabilistic models(DDPMs; (Ho et al., 2020); (Sohl-Dickstein et al., 2015); (Song et al., 2020b)), have emerged as a compelling alternative, exhibiting unmatched capabilities in generating superior samples in diverse domains, from image synthesis to molecule design (Bengio et al., 2014).
@@
Driven by these observations, our research proposes a method of integrating skip-step sampling as an optimization target into the training process. The model trained in this way can consider and adapt to the information lost during the sampling period due to skip-step sampling. Our model not only maintains efficient generation when using other accelerated sampling algorithms but also improves the quality of the generated samples, thereby achieving a balance of efficiency and performance. Specifically, during the training process, the original loss function is retained, and at the same time, a skip-step loss is introduced, measuring the difference between the model’s current step prediction and the skip-step result. This skip-step loss is combined with the original loss using a weighted mechanism. Figure 1 shows an overview of the diffusion models.
@@ 3.2 Training with Skip-Step Loss
To achieve our desired outcome, we minimize the squared difference between our targeted optimization goal and the original target, thereby effectively reducing the discrepancy between them,
After simplification following re-parameterization, we are able to derive our final loss function ,
Owing to the adjustment of certain weight values, it becomes crucial to incorporate a suitable weight in the following sections to guarantee the effective functioning of the system. At present, the stability of this loss function’s value is maintained, effectively averting any system collapse. This stability is evidenced by the results detailed in the experimental section (see the experimental in section 4).
@@ 4 Experiments
In consideration of computational resources, our experiments utilized the CIFAR10 dataset with a resolution of 32×32 and the CelebA dataset with a resolution of 64×64. For the training setup, we adopted the same architecture (He et al., 2016); (Ronneberger et al., 2015); (Kingma & Ba, 2014) as provided in the official DDIMs repository. We also ensured that all parameters were kept consistent, guaranteeing the reproducibility of our experiments. For hardware, we trained both datasets on two NVIDIA A100 GPUs, with 600K steps for CIFAR10 and 400K steps for CelebA. Model performance was evaluated based on the FID (Heusel et al., 2017); (Jolicoeur-Martineau et al., 2020). Specifically, Our evaluation method followed the DDIMs repository, wherein we sampled 50,000 images and computed the FID against real images. To further guarantee reproducibility, we fixed the random seed in our experiments, ensuring that all results are replicable. (It is noteworthy that in the experiments, the models are trained based on the same parameters and methods. The results are then obtained using different sampling algorithms. Therefore, some of the results may not be consistent with those reported in original works. This is reasonable, as the parameters trained in each report are different. Our primary comparison focuses on the diffusion models trained using our method versus those trained with the original method, specifically comparing the effects of leap-step sampling under different sampling algorithms.)
@@ 4.1 Sample Quality and Ablation Experiment
In Figure 5, we specifically focused on the training time and memory usage implications of incorporating our proposed into the model. The empirical data revealed a modest increase in training time: for instance, when integrating , the training time rose from 0.0025 seconds to 0.0031 seconds per iter on CIFAR10, and from 0.0048 seconds to 0.0056 seconds on CELEBA. Notably, this increase in training time, while measurable, remains relatively minor. Moreover, it is essential to highlight that the memory usage remained largely consistent, indicating that the incorporation of does not exert additional pressure on memory resources. This suggests that the added computational cost is minimal and, crucially, does not compromise the efficiency of the training process. Given the negligible increase in training time and stable memory usage, we believe that the trade-off is favorable, considering the significant performance improvements our method offers.
In Figures 7 and 7, we display samples from the CIFAR10 and CelebA datasets generated by models with identical numbers of sampling steps and the same sampling algorithm. These figures reveal that, under the same sampling conditions, the samples generated by S-DMs are of higher quality. Additionally, as the number of sampling steps increases, so does the quality of the samples. For instance, in CIFAR10, the images of boats and cars in the second and fourth columns demonstrate that S-DMs can produce highly accurate images in just 10 steps, whereas DDIMs yield blurrier images and require more steps for high-quality outcomes. Similarly, in CelebA, the third column presents images of a male subject, where S-DMs generate a clear hat, in contrast to the blurrier hat produced by DDIMs. Moreover, the detail in images generated by PNDMs is less pronounced compared to those by S-DMs. This difference is significant, highlighting the substantial improvement in sample quality achieved by S-DMs with the same number of sampling steps.
@@ 4.2 Interpolation and Generation Consistency
The S-DMs, drawing from the deterministic generation frameworks of DDIMs (Song et al., 2020a), PNDMs (Liu et al., 2022a), and DEIS (Zhang & Chen, 2022), also incorporate the semantic interpolation typical of implicit models like GANs (Mohamed & Lakshminarayanan, 2016). This dual approach allows S-DMs to blend structured generation with creative interpolation capabilities.
In Figure 8, the interpolation outcomes of S-DMs under various skip-step settings are displayed. The results indicate that even basic interpolation in achieves semantically rich transitions between two samples. Particularly at a ’skip’ of 50, the models generate samples with remarkable quality and detailed features. This is evidenced in images 5 to 7, where the models adeptly reproduce light and shadow effects on facial features, demonstrating the nuanced capabilities of S-DMs.
Conversely, at a of 2, the model’s behavior is more akin to the original DDIMs, leading to slightly less detailed samples. This highlights the impact of skip-step intervals on the fidelity of generated images.
Moreover, the figure shows that S-DMs maintain a consistent level of quality across different training settings, even when conditioned on the same encoding. This underscores the robustness and adaptability of S-DMs, capable of producing stable and high-quality outputs across various skip-step configurations.
@@ 5 Related Work
Denoising Diffusion Probabilistic Models (DDPMs; (Ho et al., 2020)) and Noise Conditional Score Networks (NCSNs; (Song & Ermon, 2019)) are notable for their sample quality, comparable to GANs. While DDPMs optimize a variational lower bound, NCSNs target score matching over a Parzen density estimator. Both models employ a denoising autoencoder across noise levels and use Langevin dynamics for sampling. The shared approach requires multiple iterations for optimal sample quality. Recent advancements aim to decrease DDPMs’ inference steps through SDE solvers and programming algorithms. However, challenges like the disparity between log-likelihood reduction and FID (Heusel et al., 2017), (Szegedy et al., 2016) remain in some models.
@@ 6 Conclusion and Future Work
Our findings pave a new direction for future diffusion model research. By investigating the asymmetry between the training of diffusion models and accelerated sampling algorithms, we can ensure that the trained models are better suited to accelerated sampling algorithms. This ensures that even with algorithms using fewer sampling steps, the model can still generate high-quality outputs. Perhaps this offers an effective solution to the challenges diffusion models face in balancing sampling speed and sample quality. In the future, we will continue to explore how to better integrate leap-step information to ensure greater consistency between the trained model and accelerated sampling algorithms, thereby improving sample quality. Additionally, we plan to investigate how to incorporate leap-step information into ODEs, while also exploring possibilities in non-continuous spaces.